Adaptiv fa protokoll

Egy lehetséges megoldás tehát a következő program:

T RUE ; \overline{buffer.read synchronized() ; prefix := (<>, 0) ; AKCIO ; \begin{aligned}\ atvitel \\ & \  utkozes \end{aligned} ; SKIP ; prefix := (< 0 >, 1) ; prefix.length! = 0 ; \overline{AKCIO ; \begin{aligned}\  atvitel \\ & \  utkozes \end{aligned} ; \frac{\  prefix.IP.lov = 0  /}{{prefix.IP.lov \atop := 1} ; prefix.visszalep IP () ; \  prefix.length = 0  / ; \begin{aligned}SKIP \\ & prefix.IP.lov := 1 \end{aligned}} ; prefix.IP.loext(0) ; prefix.length ; := prefix.length + 1}}

AKCIO:

\frac{\  prefix.matched(local.IP) ∧ ¬buffer.empty  /}{\begin{aligned}buffer.send synchronized() \\ & buffer.wait synchronized() \end{aligned}}